Find the inverse of the function from the ‘procedural perspective’ discussed in Example 6.1.5

f(x)=7^(x+5)-12

Answer by htmentor(1343)

f(x) = y = 7^(x+5)-12
To find the inverse, we solve the equation for x:
ln(y + 12) = (x + 5)*ln(7)
x = ln(y + 12)/ln(7) - 5
Reverse x and y:
y = ln(x + 12)/ln(7) - 5